Monetary assets=$5,000; Retirement saving assets=$12,000; Tangible assets=$43,000. Credit card debt(outbalance)=$12,000; the client didn’t calculate her detailed long-term debt amount, but he has the amount of Net worth=$34,000. How much is his total liability and how much is his long-term liability?

A. $94,000; $82,000

B. $26,000; $14;000

C. $26,000; $40,000

D. $60,000; $26,000
Answer #1

Total asset = Monetary assets=$5,000 + Retirement saving assets=$12,000+ Tangible assets=$43,000 = $60000

Total liability = Credit card debt(outbalance)=$12,000 + long term liability = x = 12000+x

Net worth = asset - liability = 34000 =60000-12000-x

x=14000 = long term liability

Total liability = 12000+x = 12000+14000=26000

so answer is option b